Startup Night 15 : UT 20140913

02:00 Cloudy at sunset, started to rain. RH98.6%

LUCI1 Calibrations

luci.20140913.NNNN.fits

Darks
  • 0001 0002 1x 4.0s dark O2DCR
  • 0003 0004 1x12.0s dark MER10

Calibrations for BRothberg
  • Scripts in /home/observer/LBTO/Rothberg/SDSS_J0122
  • 0005 0024 flat_200hk.txt. Halo2 first 10 lamp OFF, second 10 lamp ON 47k ADU in H, kinda high
  • 0025 0044 flat_200hk.txt. Changed from 20s to 10s 24k ADU in H
  • 0045 0054 1x15.0s H1+H2 OFF (manual...can't do 2 lamps in a script) See flat_210zJHK_K.txt
  • 0055 0064 1x15.0s H1+H2 ON ~12K ADU @ blue end, ~7K @ red end

03:30 RH still in limits

Note: internal disk down to 9.7GB free...clean off tonight! ...34 GB free now.

7:00 it was clearing, clouds over summit again. RH=85....90...95.

9:24 still cloudy, wet on summit. Doesn't look like it will change.

-- %USERSIG{DavidThompson - 2014-09-13}%

Comments

 
Topic revision: r3 - 13 Sep 2014, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback